An infinite line charge of uniform linear charge de sity λ--2.3 μC/m lies parallel to the y axis at x--2 m. A pont charge of 3.6 μC is located at x = 2.5 m, y = 3.5 m. Find the x component of the electric field at x 3.5 m,y-3.0 m. kN/C Enter Help
Answer #1

Electric field due to infinite line charge is given by

E=\frac{2k\lambda }{r}

2x (9 109 ) ( 2.3 x 10-6) -i--11828.57iN/C 3,5

Electric field due to point charge

\overrightarrow{E}=\frac{KQ}{|r|^{2}}\widehat{r}

r=(3.5-2.5)i + (3-3.5)j =i-0.5j

\left | \overrightarrow{r} \right |=\sqrt{1^{2}+(-0.5)^{2}}=1.118 m

= 0.89442-0.4472 1.118

(9 109)(3.6 x 10-6) E2 1.25 11591.78j N/C

The X-component of electric field

Ex=-11828.57+23183.55=11355 N/C

Ex=11.355 KN/C